In how many ways can letter of the word BALLOON be arranged so that the two Ls do not come together.
Answer by stanbon(75887)

In how many ways can letter of the word BALLOON be arranged so that the two Ls do not come together
---------------------------
Total # of arrangements: 7!/[2!*2*]
-------------------
Total # of arrangements with the LL together:
Note: this means the LL is like one letter.
# = 6!/2!
---------------------------------------
Therefore, total number with the LL NOT together:
7!/(2!*2!) - 6!/2! = 1260 - 360 = 900
========================================================
Cheers,
Stan H.